If the market demand and supply function are given as dlqd+5p=120 qs=11p- 80 then what are the market clearing price and the crossponding quantities ?
1
Expert's answer
2019-01-21T10:45:10-0500

qd + 5p = 120 or qd = 120 - 5p,

qs=11p - 80.

The market clearing price and the corresponding quantities are:

qd = qs,

120 - 5p = 11p - 80,

16p = 200,

p = 12.5,

q = 120 - 5*12.5 = 57.5.